Wednesday, August 31, 2016

SALT's back online!

What a difference 30 days can make...  The RSS was brought down on 1 August, it's been taken apart completely, extensively over-hauled & put back together.  So today it was time to get it back on top of the telescope!

Gone - almost without a trace!
Looking & working better in every way than it did when it came down, the spectrograph was wheeled out into the loading bay outside the spectrometer room.  

The lifting team getting everything set up
With RSS rigged up, released from the support frame & ready to ascend through the hatch in the floor, it was project leader (& very happy, but still anxious!) Ockert's cue to fire up the inspirational soundtrack :)

Cue the music!
The crafty scheme to pull back a day & thus restore the schedule was to lift the instrument with the detector cold.  That required keeping the cryo-cooler attached & running throughout most of the lift & installation.  It added a bit more complexity to an already daunting procedure, but it worked a treat as it allowed us to get on sky as soon as all of the cabling/fibres/hoses etc had been reconnected up there.

Cryotiger in tow this time!
An enthusiastic crowd was in place to take a last look as the spectrograph (& cryo-cooler!) paraded past the catwalk...

Quite a crowd gathered for the occasion
The structure was then rotated into place (by Ant down in the control room) & then with the landing zone within reach, the compressor was disconnected & taken back down to its igloo.

Time to disconnect the cryo-cooler
Dome crane operator Denville could then tweak the dome rotation & lower the RSS towards the tracker, to be guided down carefully by the many hands below. 

Being helped along for a smooth landing on the tracker
Timmy & Jono, perched on the edge of the world, could later crank down the chain hoists to slacken the lifting slings.

Try not to think about where Timmy was actually standing!
With all the life-support systems connected up top, the control room filled up as Fred did a coarse mirror alignment (a fine alignment was out of the question due to the high the wind).

Waiting to see whether the telescope really is back together again
The first pointing was to a spectacular target - M16, also known as the Eagle Nebula, a star-forming region in the constellation Serpens.  As usual, a picture was worth at least a thousand words!  It immediately showed that the telescope pointing's good, the image quality looks fine (given that we could not do a fine alignment of the primary because of the wind), even the guide probe mapping was still close enough to find a guide star.  

2 sec SALTICAM image of the Eagle Nebula!
Retracting the SALTICAM fold mirror in the payload then allowed the light to reach the RSS & that image looked great too...  

20 sec H-alpha image taken with RSS
A quick (5 sec) spectrum of a bright star in the field confirmed that the slitmask, grating & filter mechanisms are all working as they should too :)  There's plenty more testing & on-sky performance verification to be done, so now it's mostly over to Astro Ops to put everything through its paces & check that all's well before getting back to full science operations in a few days time (weather permitting)... 

We've blasted through August, so now it's getting Springy in Suthers
It's been an incredibly busy month up here.  Huge thanks & congratulations are in order for a superb team effort, & particularly to Ockert for doing a fantastic job of organising & leading such a challenging project!  We eag(er)ly await a photometric night to take throughput measurements!

Watch this space!

Tuesday, August 30, 2016

The payload goes back up

The big news today is that the payload & SALTICAM (in its own box) went back up on the tracker - so the telescope's coming together again!  Here are two views captured by the live-feed dome cameras...

video

video

Hopefully everything's still on track to lift the RSS tomorrow & (weather-permitting) get back on-sky in the evening!

Monday, August 29, 2016

Extreme Observing for the UCT 3rd year astronomy class!

This week we've had all the 3rd year UCT astronomy students up here in two batches, learning how to observe & reduce data.  Black belt galaxy hunters Lucas Macri & Renée Kraan-Korteweg have been pushing the 74", SpUpNIC, the students & themselves to their respective limits - obtaining spectra of ~40 galaxies per night during unseasonably good weather the past few nights.  The group visited SALT as well & thus got to see the telescope in pieces, with the RSS, SALTICAM & the payload down in the spectrometer room.  Good luck guys - hopefully you'll become regulars up here!

Time for a quick team-photo in between the arrival of group 2 & the departure of group 1
Meanwhile - the SALT team is gearing up to lift the payload tomorrow :)

Sunday, August 28, 2016

Great progress on RSS software this weekend

With the spectrograph back together & most people gone for the weekend, Ant was free to thrash the system from the software side using the new (spare) PCON machine.  A huge amount of work was done & most of the acceptance tests were successful.  A few issues remain & will be addressed over the coming days, with the hope that the RSS lift will take place on Wednesday.

To unwind after all the software testing, Ant produced yet another beautiful hill-top water colour!

Saturday, August 27, 2016

UCT science students visit the Observatory

This afternoon we got to show an enthusiastic group of about 40 first-year science students from the University of Cape Town's Extended Degree Programme (warmly led by Assistant Dean David Gammon) around the telescopes.  The students were full of energy & good questions, so will surely do well in whatever scientific fields they decide to pursue!  Tonight they return for a night tour at the visitor centre & their timing is perfect; with Saturn, Mars & the Milky Way overhead, no moon & unseasonably warm weather for what's nominally the heart of winter...

How many future astronomers can you spot in this group?

Friday, August 26, 2016

Sutherland's night sky - horizon to horizon

It's been a hugely productive week up here!  In just over four days, the RSS has gone from having lens groups & various mechanisms scattered about in different rooms - to being back together, aligned & functional.  The hexapod anti-gravity system has also been installed, so that's another major task off the list...  Midst all this, Janus still managed to find time to put together this incredible image of the Milky Way arching overhead on a moonless night.  This panorama consists of 15 individual 30 second frames (tiled 5 x 3) at ISO 6400 to minimise trailing of the stars.  & That is yet another reason why we love this place!!

The Milky Way arches overhead on a dark winter night in the Karoo (click to enlarge) - beautiful job Janus!

Thursday, August 25, 2016

Installing the hexapod anti-gravity system

Janus has produced a master(ful) document describing all the many things that need to be checked, mapped & updated once the payload & instruments go back up.  These include the telescope pointing, the auto-collimator offsets, the FIF offsets for the different HRS fibres, all the fiducials based on fold mirror positions, the various guide probes, etc etc.  Getting all of that right as soon as we get back on sky is the secret to a happy ending to the shut-down.  While better throughput & reliability will be wonderful, enthusiasm is quickly lost if the telescope's not properly functional afterwards!  Hugely important too are all the RSS tests that were done before the instrument came down & which will be repeated before it goes back up - Astro Ops is responsible for reducing those data & checking that everything's back where it belongs before the instrument leaves the spectrometer room.

Janus briefing Ant & Paul on The document while Éric reduces RSS test data
While that was happening in the control room, Johan, Nicolaas & Eben were busy up on the tracker installing the new hexapod anti-gravity system - aka the HAGS.  This looks a bit like knee-replacement surgery for the tracker, but in fact the new hexapod motors will only be installed early next year.  This temporary dislocation of the legs was necessary to allow the springs assemblies associated with the HAGS to be fitted near the apex of each pair of legs.

Johan supporting the lower section of the detached hexapod leg
This is the second of the 3 pairs of legs to be done, & the last set does not include the luxury of working from the relatively ergonomic tracker platform.  

This would be a tough enough job to do if the tracker was on the ground!
The idea is that the powerful springs push up on the tracker's non-rotating structure to unload some of the weight that would otherwise be on the ball-joints.  This will make the tracker more comfortable by relieving some of its joint pain...

Wednesday, August 24, 2016

Our Observatory: an aerial perspective!

Ant's quad-copter's still providing superb entertainment up here!  With today being one of the most perfect days imaginable - he managed to sneak in morning, afternoon & sunset flights that yielded these Absolutely Stunning shots of the plateau & surrounds...

Note the little robotic Birmingham solar telescope on the left that's already hard at work! 
SALT's shadow visits the telescope's Operations Manager's house
SALT as you will not often see it!
Pity there wasn't anyone up on the catwalk or the tracker for scale...
The view towards the south in the late afternoon
A perfect night coming up - happily SALT has not missed many of those during the well-timed shut-down
A very busy hill-top in the middle of almost nowhere

Tuesday, August 23, 2016

All-night SAMS tests

After all the attention received during the optical service, the collimator optics are now also getting some better baffling to protect the lenses & suppress stray light within the instrument.

The collimator doublet gets a lens hood
A new gear for the camera articulation was installed & extensively tested.  This is a crucial part to allow the entire RSS camera barrel to reliably swing around to the appropriate angle (twice the angle of the grating with respect to the light beam) to suit the grating setup being used.  The camera articulation angle can be up to 100 degrees for the spectrograph's highest resolution setting.  So yes, that whole section above Eben's head in the pic below can sweep out a huge arc when configuring the instrument for a given wavelength range & resolution.

Setting up the new articulation gear
An historical quirk, very much in keeping with the way astronomy works - is that the control PC for RSS is non-intuitively called PCON.  This is a hangover from the days when the instrument was called the Prime Focus Imaging Spectrograph (PFIS), & so the PFIS Control machine became, & remains, PCON...

Ant's baby - complete with lots of warnings due to various mechanisms being in pieces or undergoing testing
So much for all the daytime action...  This week we've also had our Zen Master Telescope Whisperer Fred up here to run SAMS tests with the telescope pointing to CCAS all night (weather permitting).

Captain Fred at the helm - sensing edges!
Since the telescope is offline, it can happily stare at the CCAS tower from dusk till dawn.  That makes this a perfect time to monitor the corrections being applied to the 91 mirror segments by the edge sensors running in closed-loop.  This sort of data helps Hitesh to characterise & evaluate the system's performance.

The CCAS instrument's fine microlens array produces 7 spots per segment - the spotless region in the lower right corner of the mirror is due to the shadow cast by the tracker

Monday, August 22, 2016

Wave-plates & field lens back in!

Popped in at SALT late in the afternoon, just in time to catch Jono handing the whole wave-plate mechanism to Keith, who was sitting underneath the RSS.

Serving up a wave-plate assembly
With the wave-plates re-installed, the last of the cleaned & re-filled collimator optics, the field lens group, could go back in as well.

The wave-plates fit between the main group & the field lens group of the collimator
Ockert could then check the overall alignment of the optics against the references that were established before the optics were removed.

RSS showing off all its nerves & blood vessels
That completes the RSS optical service part of the shut-down - great stuff guys!  Look forward to seeing the improved throughput of the spectrograph when we get back on-sky :)

Sunday, August 21, 2016

A few more of SALT's optics

Before all the action starts up again tomorrow, here's a chance to introduce a couple of the other important optics at SALT...  Here's a glimpse down the SALTICAM lens barrel.

Hello SALTICAM
& this beast is the ADC - the atmospheric dispersion compensator that sits just above the SAC, below the payload.  It consists of two large wedge shaped prisms, the separation of which can be changed to correct for the different extents to which the atmosphere bends red vs blue light.

The ADC getting some TLC from Janus
This is the moving baffle that lives above the ADC & blocks out stray light that's not followed the telescope's optical path.  The green reflections are due to the coatings on each of the prism surfaces. 

The carbon fibre moving baffle on top of the ADC housing

Saturday, August 20, 2016

Saturday eagle encounter

Drove up the hill for a walk this afternoon & spotted one of the eagles sitting on a rock on the SE edge of the plateau.

Enjoying the view out to the east, with Salpeterkop in the background
Strolled to within about 200 metres of it before it decided to take off & head home down the cliff to the south.

Time to go!
Those first few wing-beats are hard work...

About to disappear below the ridge
A stunning day up here - hard to believe the weather's due to pack up tomorrow, there's even snow forecast!

Friday, August 19, 2016

A well-deserved break for the team

The field lens group was cleaned & its fluid replaced this morning.  But now it's the weekend, after a very busy & productive week at SALT.  Time for the team to take a bit of a breather before finishing the RSS optical service & gearing up to re-install the payload next week.  SALT Operator Fred's up here now to run SAMS at night (weather permitting!), but otherwise it's pretty quiet for a change...

Papa black eagle on guard duty above the cliff this afternoon

Thursday, August 18, 2016

Collimator triplet oil change!

This morning the guys drained the collimator triplet & replaced the lens fluid in the 2 gaps.  Like last time, the scene was quite gruesome!  The problematic original lens fluid (which has vastly different specific gravity to the new type of fluid that's been in use since 2011) still refuses to leave the system.

Draining the coupling fluid out of the gaps in the triplet
Here's the replacement fluid slowly going in...  As the gaps fill up, most of the evidence of the old fluid disappears again.

Lens fluid transfusion in progress
The finished product looks Really Fantastic!

Sparkly triplet with surfaces cleaned & lens fluid replaced
The throughput measurements were repeated after the fluid change & showed an additional 2-3% increase blueward of 3700 A.

Before cleaning (red), after cleaning (blue) & after fluid change (green) throughput measurements for the main group
Having completed the post-fluid-change measurements, the main group was re-installed in the instrument.

Putting the main group back in this evening
The main group & field lens assemblies go into the spectrograph from below.

Jono making sure everything's lined up properly
The lens barrel is supported on a wooden frame that's carefully positioned & then slowly jacked up on a lifting platform.

Almost there - soon to be bolted into position
With that, it was time to get started on the field lens group.  A relatively simple assembly consisting of a fused quartz bottom lens (L1) & a calcium fluoride one above it (L2), with a small fluid-filled gap in between.

The calcium fluoride (L2) lens in the RSS field lens assembly
The dirty assembly was rigged up to the throughput testing system, which consists of a broadband light source, an iris to control the beam size, various combinations of lenses (to re-image the source for each group to be tested), an integrating sphere to even out the light distribution & an extremely handy little spectrograph that Ted donated from his old lab.  The optic under test is mounted on a pneumatically actuated stage that places the lenses in & out of the beam to produce the differential measurement (a huge upgrade over the old system that relied on a person sitting in the dark actuating the stage!).  The even more significant improvement is the fact that the throughput measurements over a broad wavelength range are now in the form of spectra obtained at various grating angles, rather than photometric data taken with a set of manually deployed filters.  Janus has also written a data pipeline to painlessly calibrate the spectra & display the throughput results - brilliant! 

The new throughput testing system, with the RSS field lens group installed
Here's the profile view of the field lens group in the beam.

A bit of a tight fit
This is almost the photon's view of the setup...  The light source is in the lower right corner of the image & the lab spectrograph is visible in the background on the right.  A set of "before" measurements was taken this evening so that the cleaning can be done first thing in the morning...

Field lens group in beam in the throughput testing system
While this setting up was going on in the clean room, most of the team was still busy with various tasks in the spectrometer room - only about 11 hours after they showed up here this morning!

Tell us a story Uncle Paul!
The notoriously tricky slit-mask mechanism was putting Thabelo & co. through their paces...

Feeding the monster
2.5 weeks into the shut-down, things are still on schedule & going well - great work by all concerned!